OpenStudy (anonymous):

calculate the probability of tossing a coin 15 times and getting the given number of heads. round the answer to 4 places.

OpenStudy (anonymous):

please advise what does that mean.

OpenStudy (amistre64):

what is the given number of heads?

OpenStudy (amistre64):

the answer is the binomial im sure

OpenStudy (anonymous):

the number of heads are 4, 7 and 10

OpenStudy (amistre64):

15 C #heads * .5 right?

OpenStudy (amistre64):

15.14.13.12 ----------- * 1/2 4.3.2 15.14.13.12.11.10.9 ------------------- * 1/2 7.6.5.4.3.2 15.14.13.12.11 --------------- * 1/2 5.4.3.2

OpenStudy (anonymous):

are you doing factorial? please advise

OpenStudy (amistre64):

i am.. 15 P 4 ------ * 1/2 4! 15 P 7 ------ * 1/2 7! 15 P 5 ------ * 1/2 5!

OpenStudy (zarkon):

\[_{15}C_{n}(.5)^{15}\] where n is the number of heads

OpenStudy (amistre64):

i dont think its .5^(15) tho. that looks odd to me

OpenStudy (amistre64):

or is it nCr.p.q?

OpenStudy (amistre64):

I aint got me ti83 handy to check :)

OpenStudy (zarkon):

\[P(X=x)=_{n}C_{x}p^x(1-p)^{n-x}\] follows the binomial distribution

OpenStudy (anonymous):

i have a ti83 to check....how do you put that information in?

OpenStudy (amistre64):

2nd, vars, binomialpdf(15,.5), sto, 2nd, 2 then fill in list 1 with 0,1,2,3,4,5, ... , 14,15 then stats, calc, 1-vars, L1,L2

OpenStudy (amistre64):

that last step is prolly off ... L1 tells you the number of head; and the binomial of L2 gives the probability

OpenStudy (amistre64):

1-variables stat is usually to get the mean and sd and such ...

OpenStudy (zarkon):

Let n=15 and x= # of heads \[P(X=x)=_{n}C_{x}p^x(1-p)^{n-x}\] \[=_{15}C_{x}(.5)^x(1-.5)^{15-x}\] \[=_{15}C_{x}(.5)^x(.5)^{15-x}\] \[=_{15}C_{x}(.5)^{x+15-x}\] \[=_{15}C_{x}(.5)^{15}\]

OpenStudy (anonymous):

when do press enter?

OpenStudy (zarkon):

\[_{15}C_{x}=\frac{15!}{x!(15-x)!}\]

OpenStudy (amistre64):

that way is good too :) binomialpdf(15,.5,#heads) is the direct answer without the list

OpenStudy (amistre64):

binomialpdf(15,.5,4), enter for 4 heads binomialpdf(15,.5,7), enter for 7 heads binomialpdf(15,.5,10), enter for 10 heads

OpenStudy (amistre64):

2nd, enter gets you back to what you did prior without having to go thru the distr menu again

OpenStudy (anonymous):

good answers to you both